# Debatable Intelligence: Benchmarking LLM Judges via Debate Speech Evaluation

### v1

@@ Acknowledgments
Appendix A Additional details on the data
We take the benchmarking data in this study from Slonim et al. (2021). Specifically, we use "Pipeline-set-1," a subset containing the most human-authored speeches. We removed 78 speeches that were used for vetting the annotation process, as these were designed to receive lower scores and identify poor quality annotations. We also excluded speeches from two topics — "We should increase airport racial profiling in the United States" and "We should subsidize the human mission to Mars" — since they lacked speeches from both Arg-Human1 and Arg-Human2.
@@ Appendix A Additional details on the data
Appendix B Judges implementation details
Figure 8 presents the prompt used for the zero-shot experiment, while Figure 9 presents its Chain of Thought (CoT) variation. In all experiments, we run Llama and Qwen models larger than 8B parameters using a quantized implementation (4bit). The temperature for all baselines is 0.01 and the maximal context length is 4096.
Appendix C Speech generation implementation
Figure 10 presents the prompt we use for speech generation. It requests the generation model to produce a 600-word speech supporting a specific debate topic (e.g., "We should subsidize higher education"). We use GPT-4.1 for the generation process, and create two speeches for each topic (similarly to human-authored speeches). For this step, we select a temperature of 1 and a maximal context of 1024. In case a generated speech exceeds the allowed 600 words, we iteratively remove sentences from its end until it complies with the required length.
@@ Appendix D Parsing errors
Figure 11: Number of parse errors for different prompt variations. "No-CoT" refers to the prompt described by Figure 8 and "CoT" to the prompt described by Figure 9. We note that parsing errors mostly occur with smaller models and could largely vary with the prompt. In general, results for the CoT prompt seem to be more challenging to parse.
@@
Figure 12: Results for two judge ensembles and their individual parts: Figure 12(a) shows an ensemble of relatively weak models (with $0.50.1<tau<$ 0.5) and Figure 12(b) shows a larger ensemble of more capable judges (with $tau>0.5$). Our results indicate that ensembling judges is a non-trivial enhancement technique, and might be beneficial only in certain cases.
Appendix E Judge ensembling
We inspect ensembling - aggregating the results of multiple diverse judges - as a means to improve LLM rating on this task. We examine two judge ensembles, divided according to their performance in the experiment presented in Figure 2(b). (1) Weak judges: all models with $0.50.1<tau<$ 0.5 (2) Strong judges: all models with $tau>0.5$. Figure 12 presents results for both ensembles. We observe that ensembling the set of relatively weak models (GPT-4.1-nano, Qwen-2.5-7B, and Llama-3.1-8B) offers a slight improvement of 5 points. Interestingly, ensembling models with $tau>0.5$ has no effect. These findings demonstrate that model ensembling is a complex technique which can be beneficial in some scenarios but not in others.
@@ Appendix F Additional judges scores distributions
Figure 13: Judge scores distributions. Strong LLM judges tend to give lower scores than humans (-1 signifies parsing issues)
@@
Appendix G KPA: Additional details
In Section 5.4, we describe how we use Key Point Analysis (KPA) to interpret scores assigned by one of our strongest identified judges, Llama-3.3-70B, with CoT prompting. The analysis is designed for texts with relatively short sentences, such as arguments or product reviews. To better support this analysis, we use GPT-4.1 to rephrase the raw CoT strings into paragraphs of shorter sentences. Figure 14 shows the prompt for this step. For a more robust analysis, we build the corpus using CoT reasoning from the three strongest CoT judges per model family, identified in 5.1: Qwen-32B (CoT), Llama-3.3-70B (CoT), and GPT-4.1-mini (CoT). We report results for Llama-3.3-70B (CoT) in Section 5.4. Figures 15 and 16 report similar results for GPT-4.1-mini and Qwen-32B, respectively.
@@ Appendix H Speech examples
“We should increase fuel tax. So, in this debate, we’re going to argue that the states should introduce a tax on fuels, on fuels, in order to recoup the cost of fuel. So, in the status quo, states are allowed to levy a small tax on every single product that they consume. But they can’t levy a tax on fuels, because that would be an unfair tax on people who are worse off in the world. We think that the state should instead tax fuels, and if the cost of fuel exceeds the cost of fuel, the state should tax them. So, with that in mind, we have three arguments for why we should tax the fuel. The first is that we think that this will be a fair and just solution for the climate crisis. So, right now, there’s a huge debate in the developed world about climate change. The most important question, of course, is whether we can stop global warming dead in its tracks, and the answer is no. We need to move as quickly as possible to reduce the amount of carbon that is put into the atmosphere, which will inevitably lead to global warming, which will lead to things like rising sea levels and more extreme weather. The only way to do this, therefore, is to put as many people as possible in power who can push the buttons, push the levers, and turn the lights off. But we think that, right now, there’s a huge power imbalance in the developed and developing world, and that, ultimately, the only way to make a change, to fight climate change, is to have a market that is regulated, that is controlled, that is controlled by the people who live on the ground. We think that this is the only way to do it, because the people who live on the ground are the most informed, and the people who are making the decisions are the most educated. We think that, therefore, the people who are most affected by climate change are the people on the ground, and that we need to put the people who are most affected in power. The reason for this is that, right now, the average person in the developed world is only a single vote out of a whole lot of people in the developed world. We think that, therefore, if you put a tax on the people living on the ground, that they’re more likely to take action, and that this is a one-time opportunity for people to change their behavior. The people who are in power right now are very very different from the people who are in power right now, and it’s clear that, in the future, people will have more power, and we’ll be able to use that power more efficiently. Our second argument is that this will save states money. So, right now, if states want to build up their energy infrastructure, they’re spending a lot of money to do so. And that’s a problem. We see that, for instance, in the US, that a big oil companies are using up all of the oil in the country, and that’s why the price of oil is going up. It’s a problem that, right now, the US is spending more than \$100 billion a year on oil. But, right now, we’re spending less than \$1 trillion a year on oil, which is a fraction of what we need. We think that if you tax the energy that the state uses, that that money will go much further. For instance, if you have a big nuclear power plant, and the price of nuclear power goes up, and people are using that for their own energy, that means that you’re going to be able to get more of the energy that you need for your own energy needs, and that money will go further. And, in the same way, a tax on fuel will also reduce the cost of fuel. So, ultimately, because you will be able to recoup the cost of fuel, and because this will save the state money in the long term, we think that this is a good idea. For all these reasons, we should tax fuel. Thank you.”
@@
Figures 17, 18, 19 provide examples of speeches from multiple sources: Speech-GPT2, Human expert, and GPT-4.1. All speeches discuss the same topic (“We should increase fuel tax”).

### v2

@@ Acknowledgments
Appendix A Inference Implementation Details
Figure 8: Speech scoring prompt. We give LLM judges a speech and its topic, and instruct them to rate it. The instructions closely follow the ones given to human annotators in Slonim et al. (2021).
Figure 9: Chain of Thought (CoT) speech scoring prompt. We request the LLM judge to give a short justification for the given score.
Figure 8 presents the prompt used for the zero-shot experiment, while Figure 9 presents its Chain of Thought (CoT) variation. In all experiments, we run Llama and Qwen models larger than 88B parameters using a quantized implementation (44bit). The temperature for all baselines is $0.01$ and the maximal context length is $4096$. Our experiments present results for running each judge model once.
Runs variability
We use a low temperature to promote deterministic behavior across runs. To verify this assumption, we conducted a small-scale replication with GPT-4.1, rerunning each speech three times under the same prompt. Both Tau and Kappa agreement scores remained stable, with variations no larger than $\pm{0.003}$ across runs, indicating high consistency.
Appendix B Speech Generation Implementation
@@
Figure 10 presents the prompt we use for speech generation. It requests the generation model to produce a 600-word speech supporting a specific debate topic (e.g., “We should subsidize higher education”). We use GPT-4.1 for the generation process, and create two speeches for each topic (similarly to human-authored speeches). For this step, we select a temperature of 1 and a maximal context of 1024. In case a generated speech exceeds the allowed 600 words, we iteratively remove sentences from its end until it complies with the required length.
Appendix C Additional Benchmark Details
We take the benchmarking data in this study from Slonim et al. (2021). In this section, we add information on its collection methods (§C.1), and elaborate on the partition used in this study (§C.2). For complete details, please refer to the supplementary information from their work [^2] https://www.nature.com/articles/s41586-021-03215-w#Sec14.
C.1 Original Data Collection
@@
Human-expert speeches
The dataset introduced by Slonim et al. (2021) also contains speeches produced by human experts. These were transcribed from recordings of professional debaters delivering speeches in real time (see Orbach et al. (2020); Mirkin et al. (2018)). For each topic, the dataset includes two such speeches, each recorded by a different debater.
Annotation task
Slonim et al. (2021) relied on crowd annotators to annotate speeches across sources. The annotators received the following instructions: “Imagine the following scenario. You are in the audience of a competitive debate between two opposing speakers on the specified topic. The first speaker delivers the opening speech, aiming to persuade the audience to support the topic. Please carefully read the transcript of this opening speech provided below. For each of the following statements, please indicate to what extent you agree or disagree with the statements”. They were then asked to indicate the extent to which they agreed with two statements: (1) “This speech is a good opening speech for supporting the topic”, (2) “Most arguments in this speech support the topic”. Responses were given on a five-point Likert scale, from Strongly disagree (1) to Strongly agree (5). Importantly, annotators were blind to the origin of the speech.
Annotation quality
Inter-annotator agreement for the main question (“This speech is a good opening speech for supporting the topic”) was $\kappa=0.24$, which Slonim et al. (2021) attribute to the task’s subjective nature rather than poor annotation quality. To validate reliability, they conducted three complementary checks:
  1. 1.
    Expert speeches: Speeches delivered by human debate experts received consistently higher average scores than those produced by automatic systems.
  2. 2.
    Control questions: Annotators were also asked to rate a secondary statement — “Most arguments in this speech support the topic” — which served as a control for annotator reliability. In addition, “control speeches” of intentionally low quality were included to identify inattentive annotators.
  3. 3.
    Manual review: A qualitative inspection of 20 speeches (10 high- and 10 low-scoring) confirmed that high scores corresponded to coherent, on-topic content, while low scores reflected off-topic, repetitive, or non-argumentative text.
This vetting process yielded $15$ human annotations per speech.
C.2 Data Partition Used
We use “Pipeline-set-1”, a subset containing the most human-authored speeches. We removed 78 speeches that were used for vetting the annotation process, as these were designed to receive lower scores and identify poor quality annotations. We also excluded speeches from two topics — “We should increase airport racial profiling in the United States” and “We should subsidize the human mission to Mars” — since they lacked speeches from both Arg-Human1 and Arg-Human2.
@@ Appendix D Parsing Errors
Figure 11: Number of parse errors for different prompt variations. “No-CoT” refers to the prompt described by Figure 8 and “CoT” to the prompt described by Figure 9. We note that parsing errors mostly occur with smaller models and could largely vary with the prompt. In general, results for the CoT prompt seem to be more challenging to parse.
@@
Appendix E Judge Ensembling
We inspect ensembling - aggregating the results of multiple diverse judges - as a means to improve LLM rating on this task. We examine two judge ensembles, divided according to their performance in the experiment presented in Figure 2(b). (1) Weak judges: all models with $0.1<tau<0.5$ (2) Strong judges: all models with $tau>0.5$. Figure 12 presents results for both ensembles. We observe that ensembling the set of relatively weak models (GPT-4.1-nano, Qwen-2.5-7B, and Llama-3.1-8B) offers a slight improvement of 5 points. Interestingly, ensembling models with $tau>0.5$ has no effect. These findings demonstrate that model ensembling is a complex technique which can be beneficial in some scenarios but not in others.
@@ Appendix E Judge Ensembling
Figure 12: Results for two judge ensembles and their individual parts: Figure 12(a) shows an ensemble of relatively weak models (with $0.1<tau<0.5$) and Figure 12(b) shows a larger ensemble of more capable judges (with $tau>0.5$). Our results indicate that ensembling judges is a non-trivial enhancement technique, and might be beneficial only in certain cases.
@@ Appendix F Additional Judges Scores Distributions
Figure 13: Judge scores distributions. Strong LLM judges tend to give lower scores than humans (-1 signifies parsing issues).
@@ Appendix G KPA: Additional Details
In Section 5.4, we describe how we use Key Point Analysis (KPA) to interpret scores assigned by one of our strongest identified judges, Llama-3.3-70B, with CoT prompting. The analysis is designed for texts with relatively short sentences, such as arguments or product reviews. To better support this analysis, we use GPT-4.1 to rephrase the raw CoT strings into paragraphs of shorter sentences. Figure 14 shows the prompt for this step. For a more robust analysis, we build the corpus using CoT reasoning from the three strongest CoT judges per model family, identified in 5.1: Qwen-32B (CoT), Llama-3.3-70B (CoT), and GPT-4.1-mini (CoT).
We report results for Llama-3.3-70B (CoT) in Section 5.4. Figures 15 and 16 report similar results for GPT-4.1-mini and Qwen-32B, respectively.
@@ Appendix H Speech Examples
“We should increase fuel tax. So, in this debate, we’re going to argue that the states should introduce a tax on fuels, on fuels, in order to recoup the cost of fuel. So, in the status quo, states are allowed to levy a small tax on every single product that they consume. But they can’t levy a tax on fuels, because that would be an unfair tax on people who are worse off in the world. We think that the state should instead tax fuels, and if the cost of fuel exceeds the cost of fuel, the state should tax them. So, with that in mind, we have three arguments for why we should tax the fuel. The first is that we think that this will be a fair and just solution for the climate crisis. So, right now, there’s a huge debate in the developed world about climate change. The most important question, of course, is whether we can stop global warming dead in its tracks, and the answer is no. We need to move as quickly as possible to reduce the amount of carbon that is put into the atmosphere, which will inevitably lead to global warming, which will lead to things like rising sea levels and more extreme weather. The only way to do this, therefore, is to put as many people as possible in power who can push the buttons, push the levers, and turn the lights off. But we think that, right now, there’s a huge power imbalance in the developed and developing world, and that, ultimately, the only way to make a change, to fight climate change, is to have a market that is regulated, that is controlled, that is controlled by the people who live on the ground. We think that this is the only way to do it, because the people who live on the ground are the most informed, and the people who are making the decisions are the most educated. We think that, therefore, the people who are most affected by climate change are the people on the ground, and that we need to put the people who are most affected in power. The reason for this is that, right now, the average person in the developed world is only a single vote out of a whole lot of people in the developed world. We think that, therefore, if you put a tax on the people living on the ground, that they’re more likely to take action, and that this is a one-time opportunity for people to change their behavior. The people who are in power right now are very very different from the people who are in power right now, and it’s clear that, in the future, people will have more power, and we’$l$ be able to use that power more efficiently. Our second argument is that this will save states money. So, right now, if states want to build up their energy infrastructure, they’re spending a lot of money to do so. And that’s a problem. We see that, for instance, in the US, that a big oil companies are using up all of the oil in the country, and that’s why the price of oil is going up. It’s a problem that, right now, the US is spending more than \$100 billion a year on oil. But, right now, we’re spending less than \$1 trillion a year on oil, which is a fraction of what we need. We think that if you tax the energy that the state uses, that that money will go much further. For instance, if you have a big nuclear power plant, and the price of nuclear power goes up, and people are using that for their own energy, that means that you’re going to be able to get more of the energy that you need for your own energy needs, and that money will go further. And, in the same way, a tax on fuel will also reduce the cost of fuel. So, ultimately, because you will be able to recoup the cost of fuel, and because this will save the state money in the long term, we think that this is a good idea. For all these reasons, we should tax fuel. Thank you.”
@@
Figures 17, 18, 19 provide examples of speeches from different sources: Speech-GPT2, Human expert, and GPT-4.1. All speeches discuss the same topic (“We should increase fuel tax”).